# Why We Build Local Large Language Models: An Observational Analysis from 35 Japanese and Multilingual LLMs

### v1

@@ Abstract
Why do we build local large language models (LLMs)? What should a local LLM learn from the target language? Which abilities can be transferred from other languages? Do language-specific scaling laws exist? To explore these research questions, we evaluated 35 Japanese, English, and multilingual LLMs on 19 evaluation benchmarks for Japanese and English, taking Japanese as a local language. Adopting an observational approach, we analyzed correlations of benchmark scores, and conducted principal component analysis (PCA) on the scores to derive ability factors of local LLMs.
We found that training on English text can improve the scores of academic subjects in Japanese (JMMLU). In addition, it is unnecessary to specifically train on Japanese text to enhance abilities for solving Japanese code generation, arithmetic reasoning, commonsense, and reading comprehension tasks. In contrast, training on Japanese text could improve question-answering tasks about Japanese knowledge and English-Japanese translation, which indicates that abilities for solving these two tasks can be regarded as Japanese abilities for LLMs. Furthermore, we confirmed that the Japanese abilities scale with the computational budget for Japanese text.
Why We Build Local Large Language Models:
An Observational Analysis from 35 Japanese and Multilingual LLMs
Koshiro Saito1, Sakae Mizuki2,1 Masanari Ohi1, Taishi Nakamura1, Taihei Shiotani1, Koki Maeda1, Youmi Ma1, Kakeru Hattori1, Kazuki Fujii1, Takumi Okamoto1, Shigeki Ishida1, Hiroya Takamura2, Rio Yokota1, Naoaki Okazaki1, 1Institute of Science Tokyo 2National Institute of Advanced Industrial Science and Technology
@@ 1 Introduction
Major large language models (LLMs) are English-centric (English LLMs hereafter), e.g., Meta Llama 3 (Dubey et al., 2024), Mistral (Jiang et al., 2023), and Phi-3 (Abdin et al., 2024), due to the dominance of English on the internet and the global economy, which results in a limited focus on non-English languages. Several companies and research institutes have been actively developing LLMs that perform well on non-English texts (local LLMs hereafter), e.g., Bllossom (Choi et al., 2024), Chinese-LLaMA (Cui et al., 2024) and openCabrita (Larcher et al., 2023), driven by various motivations. These include advancing research and development in natural language processing, mitigating security risks associated with relying on a limited number of foreign companies, and promoting responsible artificial intelligence for their community.
Recently, dozens of Japanese-centric LLMs (Japanese LLMs hereafter) have been developed in Japan, such as Sarashina2111https://huggingface.co/sbintuitions/sarashina2-7b, Llama 3.1 Swallow [^2] https://swallow-llm.github.io/llama3-swallow.en.html, and LLM-jp (LLM-jp et al., 2024). However, the advantages of training LLMs on non-English text, such as Japanese, remain underexplored. While LLMs have demonstrated high multilingual abilities, such as arithmetic reasoning (Shi et al., 2023) and machine translation (Briakou et al., 2023), Zhang et al. (2023) reported that they performed poorly on non-English commonsense reasoning. Although Fujii et al. (2024) reported that training on Japanese text improved question-answering (QA) tasks, the contribution of Japanese training data on each task has not been investigated.
A straightforward approach for analyzing the impact of Japanese training data in LLMs is to conduct ablation studies; more specifically, we prepare training data by changing the size and mix of Japanese data sets, train an LLM on the data, and measure its performance. However, this approach is inefficient due to the significant computational resources required. Even if such studies were conducted, it would remain unclear whether the findings could be generalized beyond specific design choices, such as variations in training data, the numbers of model parameters, or pre-training methods (from scratch or continual pre-training).
Instead of cost-intensive ablation studies, this paper adopts an observational approach (Ruan et al., 2024) for Japanese LLMs, leveraging the exceptionally active development of Japanese LLMs among non-English initiatives. Specifically, we evaluate publicly available 35 Japanese, English, and multilingual LLMs representing a variety of design choices. We also use 19 comprehensive evaluation benchmarks, including Japanese translations of English benchmarks, where the same tasks are evaluated in both languages. Our goal is to derive insights that are generalizable (unrestricted to specific design choices) through a quantitative analysis focusing on the following points.
@@
First, to explore multilinguality, we analyzed score correlations across 19 task benchmarks for 35 LLMs, and conducted Principal Component Analysis (PCA) to represent the performance in a low-dimensional ability space (Ruan et al., 2024). We found that tasks such as academic subjects, code generation, and arithmetic reasoning exhibited strong cross-lingual correlations on their scores and were associated with the same ability factors across languages. This indicates strong multilingual transferability, suggesting that training in English text would also improve performance in Japanese for these tasks. Conversely, tasks such as QA about Japanese knowledge and English-Japanese translation exhibited weak correlations with other English tasks and were strongly associated with an independent ability factor, which indicates language-specific abilities.
@@
Second, to examine language-specific scaling laws, we defined the language-specific computational budget as the product of the number of parameters and training tokens for each language (Hoffmann et al., 2022), and analyzed the log-linear relationship between these budgets and the ability factors obtained by PCA. We found that the English computational budget showed a strong correlation with the general ability factor but a weak correlation with the Japanese-specific ability factor. In contrast, the Japanese computational budget showed a strong correlation with the Japanese ability factor, suggesting that tasks such as QA about Japanese knowledge and English-Japanese translation scale with the amount of Japanese text and are difficult to learn solely on English texts.
@@ 2.1 Correlations between Tasks and Ability Factors
2.2 Effects of Training on Non-English Text
There is a growing number of studies examining both the promising and disappointing impacts of training local LLMs on target language data.
On the promising side, continual pre-training (continued pre-training) of strong English LLMs on non-English languages such as Chinese (Cui et al., 2024), Korean (Choi et al., 2024), Portuguese (Larcher et al., 2023), and Thai (Pipatanakul et al., 2023) has reported improvements on a variety of tasks in the target languages, including commonsense reasoning, reading comprehension, question answering, and academic subjects. Tejaswi et al. (2024) conducted systematic experiments under continual pre-training settings and reported that the effectiveness varies across different base LLMs.
On the disappointing side, Berend (2022) reported that multilingual training does not always improve performance due to the curse of multilingualty (Conneau et al., 2020). In addition, Holmström et al. (2023) reported that a Swedish LLM trained from scratch performed poorly compared to GPT-3, highlighting the difficulty of outperforming strong multilingual LLMs. Furthermore, English and multilingual LLMs reportedly show strong multilingual abilities on tasks such as arithmetic and commonsense reasoning (Shi et al., 2023) through cross-language generalization (Zhang et al., 2023). These findings suggest that the benefits of training on non-English text might be limited or, at the very least, task-dependent.
Despite these debates, there has not yet been a comprehensive and cross-lingual benchmarking using a wide variety of LLM families to assess the effect of training on non-English text.
@@ 3.1 Models
To obtain generalizable insights, we evaluated publicly available 35 Japanese, English, and Multilingual LLMs (see Table 1 in Appendix A.1 for the complete list), which represent diverse design choices, including training data, the number of model parameters, and pre-training approach. The evaluated models include: English LLMs (e.g., Llama 3 (Dubey et al., 2024), Mistral (Jiang et al., 2023), and Mixtral (Jiang et al., 2024)); Japanese LLMs continually pre-trained from English base LLMs on 18–175 billion tokens of Japanese text (e.g., Llama 3 Swallow (Fujii et al., 2024) and Llama 3 Youko (Sawada et al., 2024)); Japanese LLMs pre-trained primarily on 130–1,050 billion tokens of Japanese text from scratch (e.g., LLM-jp (LLM-jp et al., 2024) and Sarashina2; and multilingual LLMs pre-trained on multilingual data including Japanese (e.g., C4AI Command-R [^3] https://huggingface.co/CohereForAI/c4ai-command-r-v01 and Qwen2 (Yang et al., 2024)). Notably, all the English LLM families that served as base models for the continually pre-trained Japanese LLMs were evaluated as well. We focused on base models and did not evaluate instruction-tuned models to examine the effect of pre-training and avoid the confounding effects of task-oriented instruction tuning.
To estimate the computational budget for each model, we collected data on the number of model parameters and the number of training tokens for Japanese, English, and total across all languages from official sources such as technical reports, press-release documents, and model cards. Refer to Appendix A.3 for details. For a continually pre-trained model, we calculated the total number of training tokens by summing the tokens used in both initial and continual pre-training stages.
@@ 3.2 Evaluation Tasks and Benchmarks
We evaluated the models using 19 evaluation benchmarks in both Japanese and English [^4] The evaluation scores for each model are publicly available on Zenodo, licensed under CC BY 4.0, at https://doi.org/10.5281/zenodo.13959137. , which is listed in Table 2 of Appendix A.2. These tasks were selected from the perspective of cross-lingual benchmarking and comprehensiveness for general-purpose LLMs. The evaluation was conducted using zero-shot or few-shot in-context learning settings depending on tasks. Refer to Appendix A.2 for details.
We employed some Japanese benchmarks corresponding to their English counterparts for cross-lingual benchmarking: code generation (JHumanEval (Sato et al., 2024) vs. HumanEval (Chen et al., 2021)), commonsense (JCommonsenseQA (Kurihara et al., 2022) vs. XWINO (Tikhonov and Ryabinin, 2021) and HellaSwag (Zellers et al., 2019)), arithmetic reasoning (MGSM (Shi et al., 2023) vs. GSM8K (Cobbe et al., 2021)), encyclopedic knowledge-based QA (JEMHopQA (Ishii et al., 2023) and NIILC (Sekine, 2003) vs. TriviaQA (Joshi et al., 2017)), reading comprehension (JSQuAD (Kurihara et al., 2022) vs. SQuAD2 (Rajpurkar et al., 2018)), and academic subjects (JMMLU (Yin et al., 2024) vs. MMLU (Hendrycks et al., 2021)). Notably, MGSM, JMMLU, and JHumanEval are translations of GSM8K, MMLU, and HumanEval, respectively. Cross-lingual correlations between these benchmarks provide insights into the multilinguality and language specificity of each task. It is also worth noting that JEMHopQA and NIILC are developed based on Japanese Wikipedia and include instances that assess knowledge specific to Japanese culture, such as history, geography, notable figures and society, making them suitable for evaluating how much LLMs acquire knowledge about Japan.
For comprehensiveness, inspired by the natural language processing taxonomy (Chang et al., 2024; Guo et al., 2023) and to capture as many ability factors as possible, we included additional task benchmarks beyond cross-lingual benchmarks. Specifically, we employed Japanese automatic summarization (XL-Sum (Hasan et al., 2021)), machine translation between English and Japanese (WMT20-en-ja and ja-en (Barrault et al., 2020)), English question answering (OpenBookQA (Mihaylov et al., 2018)), and logical reasoning (Big-Bench-Hard (Suzgun et al., 2023)). Because we posit that local LLMs serve as foundational models for the target language, our evaluation focused on fundamental knowledge and skills rather than domain-specific tasks (e.g., question answering in financial or medical domains). Furthermore, we excluded safety and bias-related tasks, as these should be addressed in the post-training stage.
@@ 4 Experimental Results
Based on the experimental setting explained in the previous section, we obtained a matrix of benchmark scores $X\in\mathbb{R}^{M\times D}$, where $M$ and $D$ are the numbers of LLMs and benchmarks, respectively ($M=35$ and $D=19$ in this study) and an element $X_{i,j}$ presents the score of the LLM $i$ on the benchmark $j$. In this section, we analyze the matrix $X$ to unveil the difference in training strategies for Japanese LLMs (§ 4.1), similarity of benchmarks in terms of LLM performance (§ 4.2), ability factors of LLMs (§ 4.3). We then confirm that our findings about ability factors are aligned with the scaling laws (§ 4.4) and generalizable to different training strategies (§ 4.5).
@@ 4.1 Comparison of Benchmark Scores by Pre-trained Languages
Figure 1 presents a bubble chart showing the benchmark score distributions grouped by the primary language of the LLMs: Japanese continually pre-trained (light blue), Japanese trained from scratch (green), English (red), and Multilingual (gray). The variable $n$ in each group represents the number of models included.
On overall, it is evident that LLMs with larger parameters tend to achieve higher scores in each group. When comparing benchmark scores for smaller models, there is a clear tendency for LLMs continually pre-trained on Japanese text (the green bubbles) to outperform English LLMs (the red bubbles) on Japanese benchmarks (shown in blue) except JHumanEval and MGSM. This indicates the effectiveness of continual pre-training on Japanese text. The advantage is particularly evident in tasks such as Japanese QA (NIILC) and English-Japanese translation (WMT20-en-ja). Refer to Appendix B for detailed discussion. Considering their smaller number of parameters, Japanese LLMs trained from scratch (green bubbles) achieve competitive scores on most Japanese benchmarks, with the exceptions of the arithmetic reasoning (MGSM) and the code-generation (JHumanEval).

### v2

@@ Abstract
Why do we build local large language models (LLMs)? What should a local LLM learn from the target language? Which abilities can be transferred from other languages? Do language-specific scaling laws exist? To explore these research questions, we evaluated 35 Japanese, English, and multilingual LLMs on 19 evaluation benchmarks for Japanese and English, taking Japanese as a local language. Adopting an observational approach, we analyzed correlations of benchmark scores, and conducted principal component analysis (PCA) to derive ability factors. We found that if LLMs perform well in English on tasks like academic subjects, code generation, arithmetic reasoning, commonsense, and reading comprehension, they also perform well on the same tasks in Japanese. This indicates it is not necessary to specifically train on Japanese text to enhance abilities for solving these tasks. In contrast, training on Japanese text improves question-answering tasks about Japanese knowledge and English-Japanese translation, which indicates that abilities for solving these two tasks can be regarded as Japanese abilities. Furthermore, we confirmed that the Japanese abilities scale with the computational budget for Japanese text. Taken together, our findings offer generalizable insights into which tasks benefit from local-language data and what we can expect when building local LLMs.
@@ 1 Introduction
Major large language models (LLMs) are English-centric (English LLMs hereafter), e.g., Meta Llama 3 (Dubey et al., 2024), Mistral (Jiang et al., 2023), and Phi-3 (Abdin et al., 2024), due to the dominance of English on the internet and the global economy, which results in a limited focus on non-English languages. Several companies and research institutes have been actively developing LLMs targeting non-English languages (local LLMs hereafter), e.g., Bllossom (Choi et al., 2024), Chinese-LLaMA (Cui et al., 2024) and openCabrita (Larcher et al., 2023), driven by various motivations. These include advancing research and development in multilingual NLP, mitigating security risks associated with relying on a limited number of foreign companies, and promoting responsible artificial intelligence for their community.
However, the advantages of training LLMs on non-English text remain underexplored–particularly regarding the unique skills or knowledge such LLMs might gain compared to English-centric or Multilingual LLMs. On the one hand, LLMs have demonstrated high multilingual abilities, such as arithmetic reasoning (Shi et al., 2023) and machine translation (Briakou et al., 2023), which casts doubt on the advantage of training on non-English text. On the other hand, training on non-English text has been reported to bring stronger cultural and regional knowledge of the target language (Romanou et al., 2025), although there are mixed findings for other tasks such as commonsense reasoning and reading comprehension (Cui et al., 2024; Choi et al., 2024; Larcher et al., 2023). These two perspectives–multilinguality versus language specificity–suggest that the effectiveness of training on non-English text is inherently task dependent. Indeed, demonstrating an advantage of training on non-English text remains not straightforward. Numerious studies have built non-English LLMs from scratch (Holmström et al., 2023) or via continual pre-training (CPT) over English LLMs (Cui et al., 2024; Choi et al., 2024; Larcher et al., 2023), but their task-specific results are often mixed or contradictly, raising doubts about generalizability (§ 2.1). Because LLM performances depends on several design choices–such as training from scratch or via CPT, which base model is selected for CPT (Tejaswi et al., 2024a), and how the training data is curated (Penedo et al., 2024; Li et al., 2024)–it is difficult to isolate performance gains specifically attributable to training on non-English text. Given its huge impact, thorough investigation and convincing insights into the advantages of local LLMs are valuable.
To explore what unique skills or knowledge may emerge as the natural consequence of the training on non-English text, we adopt an observational approach (Ruan et al., 2024) for Japanese-centric LLMs (Japanese LLMs hereafter), leveraging the exceptionally active development in Japan (e.g., Llama 3.1 Swallow [^1] https://swallow-llm.github.io/llama3-swallow.en.html and LLM-jp (LLM-jp et al., 2024)) among non-English initiatives. Specifically, we evaluate 35 publicly available Japanese, English, and multilingual LLMs representing a variety of design choices. We also use 19 comprehensive evaluation benchmarks covering knowledge-based QA, academic subjects, reading comprehension, and more, tasked in Japanese and English. These also includes paired Japanese and English benchmarks so that we can compare the task performance across both languages. Our goal is to derive generalizable insights (i.e., insights that are robust to design choices) by conducting a quantitative analysis.
@@
First, to explore multilinguality versus language specificity, we analyzed score correlations across 19 task benchmarks for 35 LLMs, and applied Principal Component Analysis (PCA) to represent the performance in a low-dimensional ability space (Ruan et al., 2024). We found that tasks such as academic subjects, code generation, and arithmetic reasoning exhibited strong cross-lingual correlations on their scores and were associated with the same ability factors across languages. This indicates strong multilingual transferability, suggesting that training in English text would also improve performance on these tasks when tested in Japanese. Conversely, tasks such as QA about Japanese cultural knowledge and English-Japanese translation exhibited weak correlations with other tasks and were strongly associated with an independent ability factor, indicating language-specific abilities.
@@
Second, to investigate the language-specific abilities attributed to training on Japanese text, we examined language-specific scaling laws. Specifically, we defined the language-specific computational budget as the product of the number of parameters and training tokens for each language (Hoffmann et al., 2022), and analyzed the log-linear relationship between these budgets and the ability factors obtained by PCA. We found that the English computational budget showed a strong correlation with the general ability factor but a weak correlation with the Japanese-specific ability factor. In contrast, the Japanese computational budget showed a strong correlation with the Japanese ability factor, suggesting that enhancement of Japanese knowledge and English-Japanese translation skills arise from training on Japanese text itself beyond particular design choice. These knowledge and skill scale with the amount of Japanese training text and are difficult to acquire solely from English text.
@@ 2 Related Work
2.1 Effects of Training on Non-English Text
There is a growing number of studies examining the impacts of training local LLMs on target language data: Chinese (Zhao et al., 2024; Cui et al., 2024), Turkish (Toraman, 2024), Portuguese (Larcher et al., 2023), Swedish (Holmström et al., 2023), and Finnish (Luukkonen et al., 2023). Some studies consistently reported gains in reading comprehension (Etxaniz et al., 2024b; Fujii et al., 2024; Dou et al., 2024; Joshi et al., 2025; Vo et al., 2024; Larcher et al., 2023), commonsense reasoning (Etxaniz et al., 2024b; Fujii et al., 2024; Phasook et al., 2024; Dou et al., 2024; Joshi et al., 2025; Vo et al., 2024; Choi et al., 2024; Tejaswi et al., 2024b), and local knowledge QA (Etxaniz et al., 2024b; Fujii et al., 2024; Joshi et al., 2025; Etxaniz et al., 2024a). However, following our survey of 15 previous reports on non-English LLMs (see Table 1 in § A), the evidence remains fragmented for two reasons: 1) Sparse coverage of task types: Prior works evaluated only a small set of benchmarks (an average of 2.5). In particular, machine-translation and coding tasks appear in just 2 and 0 out of 15 studies, respectively. 2) Contradictory results: Some studies drew (self-)contradictory conclusions: e.g., for mathematical reasoning, Etxaniz et al. (2024b) reported positive+neutral effects, whereas Pipatanakul et al. (2023) noted negative+neutral effects; for academic subject, both of Phasook et al. (2024) and Dou et al. (2024) documented positive+neutral effects; and, for summarization, Fujii et al. (2024) observed a negative effect, whereas Joshi et al. (2025) and Tejaswi et al. (2024b) found a positive effect.
2.2 Multilinguality vs Language-specificity
Training on non-English corpora sometimes involve using multilingual corpora. Berend (2022) and Chang et al. (2024a) reported that multilingual training does not always improve performance due to the curse of multilingualty (Conneau et al., 2020). Furthermore, English and multilingual LLMs reportedly show strong multilingual abilities on tasks such as arithmetic and commonsense reasoning (Shi et al., 2023) through cross-language generalization (Zhang et al., 2023). These findings suggest that the benefits of training on non-English text might be limited or task-dependent.
@@ 3.1 Models
To obtain generalizable insights, we evaluated publicly available 35 Japanese, English, and Multilingual LLMs (see Table 2 in Appendix B.1 for the complete list), which represent diverse design choices, including training data, the number of model parameters, and pre-training approach. The evaluated models include: English LLMs (e.g., Llama 3 (Dubey et al., 2024), Mistral (Jiang et al., 2023), and Mixtral (Jiang et al., 2024)); Japanese LLMs continually pre-trained from English base LLMs on 18–175 billion tokens of Japanese text (e.g., Llama 3 Swallow (Fujii et al., 2024) and Llama 3 Youko (Sawada et al., 2024)); Japanese LLMs pre-trained primarily on 130–1,050 billion tokens of Japanese text from scratch (e.g., LLM-jp (LLM-jp et al., 2024) and Sarashina2; and multilingual LLMs pre-trained on multilingual data including Japanese (e.g., C4AI Command-R [^2] https://huggingface.co/CohereForAI/c4ai-command-r-v01 and Qwen2 (Yang et al., 2024)). Notably, all the English LLM families that served as base models for the continually pre-trained Japanese LLMs were evaluated as well. We focused on base models and did not evaluate instruction-tuned models to examine the effect of pre-training and avoid the confounding effects of task-oriented instruction tuning.
To estimate the computational budget for each model, we collected data on the number of model parameters and the number of training tokens for Japanese, English, and total across all languages from official sources such as technical reports, press-release documents, and model cards. Refer to Appendix B.3 for details. For a continually pre-trained model, we calculated the total number of training tokens by summing the tokens used in both initial and continual pre-training stages.
@@ 3.2 Evaluation Tasks and Benchmarks
We evaluated the models using 19 evaluation benchmarks in both Japanese and English [^3] The evaluation scores for each model are publicly available on Zenodo, licensed under CC BY 4.0, at https://doi.org/10.5281/zenodo.13160661. , which is listed in Table 3 of Appendix B.2. These tasks were selected from the perspective of cross-lingual benchmarking and comprehensiveness for general-purpose LLMs. The evaluation was conducted using zero-shot or few-shot in-context learning settings depending on tasks. Refer to Appendix B.2 for details.
We employed some Japanese benchmarks corresponding to their English counterparts for cross-lingual benchmarking: code generation (JHumanEval (Sato et al., 2024) vs. HumanEval (Chen et al., 2021)), commonsense (JCommonsenseQA (Kurihara et al., 2022) vs. XWINO (Tikhonov & Ryabinin, 2021) and HellaSwag (Zellers et al., 2019)), arithmetic reasoning (MGSM (Shi et al., 2023) vs. GSM8K (Cobbe et al., 2021)), encyclopedic knowledge-based QA (JEMHopQA (Ishii et al., 2023) and NIILC (Sekine, 2003) vs. TriviaQA (Joshi et al., 2017)), reading comprehension (JSQuAD (Kurihara et al., 2022) vs. SQuAD2 (Rajpurkar et al., 2018)), and academic subjects (JMMLU (Yin et al., 2024) vs. MMLU (Hendrycks et al., 2021)). Notably, MGSM, JMMLU, and JHumanEval are translations of GSM8K, MMLU, and HumanEval, respectively. Cross-lingual correlations between these benchmarks provide insights into the multilinguality and language specificity of each task. It is also worth noting that JEMHopQA and NIILC are developed based on Japanese Wikipedia and include instances that assess knowledge specific to Japanese culture, such as history, geography, notable figures and society, making them suitable for evaluating how much LLMs acquire knowledge about Japan.
For comprehensiveness, inspired by the natural language processing taxonomy (Chang et al., 2024b; Guo et al., 2023) and to capture as many ability factors as possible, we included additional task benchmarks beyond cross-lingual benchmarks. Specifically, we employed Japanese automatic summarization (XL-Sum (Hasan et al., 2021)), machine translation between English and Japanese (WMT20-en-ja and ja-en (Barrault et al., 2020)), English question answering (OpenBookQA (Mihaylov et al., 2018)), and logical reasoning (Big-Bench-Hard (Suzgun et al., 2023)). Because we posit that local LLMs serve as foundational models for the target language, our evaluation focused on fundamental knowledge and skills rather than domain-specific tasks (e.g., question answering in financial or medical domains). Furthermore, we excluded safety and bias-related tasks, as these should be addressed in the post-training stage.
@@ 4 Experimental Results
[…]
@@ 4.1 Comparison of Benchmark Scores by Pre-trained Languages
Figure 1 presents a bubble chart showing the benchmark score distributions grouped by the primary language of the LLMs: Japanese continually pre-trained (green), Japanese trained from scratch (light blue), English (red), and Multilingual (gray). The variable $n$ in each group represents the number of models included.
On overall, it is evident that LLMs with larger parameters tend to achieve higher scores in each group. When comparing benchmark scores for smaller models, there is a clear tendency for LLMs continually pre-trained on Japanese text (the green bubbles) to outperform English LLMs (the red bubbles) on Japanese benchmarks (shown in blue) except JHumanEval and MGSM. This indicates the effectiveness of continual pre-training on Japanese text. The advantage is particularly evident in tasks such as Japanese QA (NIILC) and English-Japanese translation (WMT20-en-ja). Refer to Appendix C for detailed discussion. Similarly, Japanese LLMs trained from scratch (the light blue bubbles), despite having relatively few parameters, achieve competitive scores on most Japanese benchmarks, with the exceptions of the arithmetic reasoning (MGSM) and the code-generation (JHumanEval).